OpenStudy (anonymous):

Given f(x) = 10- 16/x, find all c in the interval [-2,8] that satisfies the Mean Value Theorem.

OpenStudy (anonymous):

Begin by taking the derivative ...

OpenStudy (anonymous):

16/c^2 ?

OpenStudy (anonymous):

Now find the slope using (-2, f(-2)) and (8, f(8))

OpenStudy (anonymous):

i plug the -2 and 8 into the original equation for the top part correct?

OpenStudy (anonymous):

Yes, plug both -2 and 8 into the equation.

OpenStudy (anonymous):

For f(8) = 10 - 16/8 = 8

OpenStudy (anonymous):

is the slope one?

OpenStudy (anonymous):

Where to show my work? f(-2) = 10 + 8 = 18 Slope of (-2, 18) and (8,8) is (18-8)/(-2 - 8) = -1 correct?

OpenStudy (anonymous):

oh ok yeah i was adding the bottom ones

OpenStudy (anonymous):

Set your first answer equal to -1 and solve for c ...

OpenStudy (anonymous):

wait which one sorry

OpenStudy (anonymous):

16/c^2 = -1

OpenStudy (anonymous):

We have the slope of the endpoints equal to your derivative

OpenStudy (anonymous):

got it thanks